OpenStudy (ms_sp):

Find three consecutive odd integers such that their sum is five more than four times the largest integer.

OpenStudy (lannyxx):

Yeah convert english to math is what is being taught here. three consecutive odd numbers means they come in order, (3,5,7) so let's say the first number is o we have o, o+2, and o+4 (if o = 3, then 3+2 = 5, 3+4 = 7, it's every other number) "their sum is fire more than four times the largest integer" so, what's the largest integer??? o+4 right? their sum is o + o+2 + o+4 = 3o+6 so (3o+6) is five more than four times (o+4) (3o+6) = 5 + 4(o+4) solve for o

OpenStudy (welshfella):

you can write the 3 odd integers as 2x+1 , 2x + 3 and 2x + 5

OpenStudy (lannyxx):

oh I see, so instead of writing o, which gives the possibility that o can be even you write 2x+1 to make sure it's odd! good catch

OpenStudy (welshfella):

yes

OpenStudy (ms_sp):

In the book, I am getting the answers of -15, -13, and -11 clearly confused.

OpenStudy (welshfella):

2x + 1 + 2x + 3 + 2x + 5 = 4(2x + 5)+ 5

OpenStudy (welshfella):

6x + 9 =8x + 25 -16 = 2x x = -8 so first number = 2(-8) + 1 = -15

OpenStudy (ms_sp):

Thank you very much @welshfella , and @LannyXX!

OpenStudy (lannyxx):

yeah!